i am having a lot of trouble with my relations and functions assignment i have made the first two problems a succes after various tries, but the third is a little trickier will you please help... and show steps.(f(x)= (-11))
3. F(-11)= (-4+7(-11))/(5+5(-11)^2+5(-11))

Answer by nerdybill(7384)

Apply the "order of operations" -- PEMDAS (Please Excuse My Dear Aunt Sally) refers to:
Parentheses
Exponents
Multiplication/division -- right to left
Addition/subtraction -- right to left
.
F(-11)= (-4+7(-11))/(5+5(-11)^2+5(-11))
F(-11)= (-4+7(-11))/(5+5(121)+5(-11))
F(-11)= (-4-77)/(5+5(121)+5(-11))
F(-11)= (-4-77)/(5+605+5(-11))
F(-11)= (-4-77)/(5+605-55)
F(-11)= -81/(5+605-55)
F(-11)= -81/(610-55)
F(-11)= -81/555
F(-11)= -27/185 (this would be the simplified fraction form)
.
F(-11)= -.14595 (or, this would be the approx decimal form)